# Model-free Forecasting of Rogue Waves using Reservoir Computing

### v1

@@ 4 Maximal Intensity Breather
Figure 7: A fifth-order maximum intensity breather obtained from [53] as testing data.
We use the same reservoir, RC_AB, as in Section 3 without retraining. This means that we use the generated $W_{x},W_{u}$ along with the learned $W_{o}$ to generate hidden states $\mathbf{x}^{(k)}$ from the input testing data shown in Fig. 7. We discard $N_{w}$ transient points before calculating one-step predictions, as presented in Fig. 8. The NRMSE of the one-step prediction is displayed in panel (a), with the error axis shown on a logarithmic scale. The figure shows that the one-step error is generally insignificant, except near the peak of the breather. Even though the multiple lobes are not present in the RC training data, the RC’s one-step prediction profile is remarkably similar to the expected result, as observed in panels (b) and (c). Furthermore, although the RC prediction slightly underestimates the highest peak, as shown in panel (d), the RC successfully captures the overall shape of the highest peak breather.
Figure 8: RC_AB can accurately forecast the dynamics of the maximum intensity breather in Fig. 7 using one-step prediction, even though the RC was trained using data in Fig. 2. (a) The NRMSE of the RC prediction is shown on a logarithmic scale. (b)-(d) Snapshots comparing RC prediction and maximal intensity breather data at $t=4.75,5.75,6.25$.
Figure 9 shows the prediction horizon of the maximal intensity breather testing data, with a legend similar to Fig. 4. Overall, the prediction horizon for the maximum intensity breather is shorter than in Section 2.3. Despite the shorter prediction horizon, our RC can forecast the occurrence of extreme rogue waves with a lead time of 0.5 time units.
The maximal intensity breather represents a path within the homoclinic orbit. It is important to note that the RC model can provide reasonable one-step predictions, even though the original training data do not directly traverse this trajectory. However, due to the inherent sensitivity of the homoclinic orbit — where any infinitesimal deviation leads to exponential divergence — RC struggles with long-term autonomous prediction. The rapid divergence of trajectories near the homoclinic orbit highlights the challenge of maintaining accurate forecasts beyond short horizons.
Figure 9: RC can perform short-term forecasting of the propagation of the maximal intensity breather. The blue dots represent prediction horizons. The red line represents the maximum of the wave envelope.
@@ 5 Random Ocean Wave Simulation
This section demonstrates the forecasting performance of RC in predicting the occurrence of rogue waves in random ocean waves. To that end, we synthesize long-crested irregular waves in deep water. The initial condition for the simulation follows the JONSWAP spectrum, with the detailed procedure given in Appendix C. The physical conditions correspond to deep water waves with a peak period $T_{p}=8$ seconds and a significant wave height $H_{s}=8$ meters. According to the linear ocean wave dispersion relation, this corresponds to a water depth of at least 50 meters and a significant wave steepness of $H_{s}/L_{p}=0.08$, where $L_{p}$ is the wavelength at the peak period. We then generate several sets of data with different random phase realizations.
@@
Figure 10: RC can capture the location of rogue waves in random ocean waves for one-step prediction. The plot shows the envelope profile at three distinct times: $t=1.92,4.78,9.63$.
Figure 11: Autonomous prediction for random ocean waves. The bottom curve shows the maximal envelope amplitude. The upper figure corresponds to the profile of autonomous prediction at times corresponding to (a), (b), and (c).
For autonomous prediction, we initiate the process 0.15 NLS time units before the rogue wave event, which corresponds to 12 seconds in physical time. Overall, both RC_Cm and RC_AB demonstrate reasonable agreement with the NLS simulation data in predicting the occurrence and location of rogue waves. However, it is noteworthy that RC_Cm provides a more accurate representation of the wave envelope’s amplitude compared to RC_AB.
@@ 6 Enhancing Autonomous Prediction
$\tilde{u}^{(k)}_{j}=(1-\delta_{j})u_{j}^{(k)}+\delta_{j}\hat{u}_{j}^{(k)},$
@@
Figure 12: Heatmap of rogue wave dynamics. The upper left panel shows the ground-truth data. The remaining panels show the autonomous RC_AB prediction with modified procedures: partial update (upper right), normalization of the solution norm (lower left), and a combination of partial update and normalization of the solution norm (lower right). The update and normalization procedures accurately capture both the amplitude and timing of the rogue wave.
@@
Overall, all approaches successfully predict the occurrence of rogue waves. However, each method exhibits distinct limitations. The update-only approach tends to underestimate the rogue wave amplitude, while the normalization approach predicts the rogue wave with a slight delay. In contrast, the combined approach accurately captures both the amplitude and timing of the rogue wave.
@@
Next, we examine the sensitivity of the update interval in the combined approach by considering three assimilation intervals: 30, 70, and 90 time steps. The top left panel of the figure shows the peak profile of the forecasted rogue wave, demonstrating close agreement in both shape and amplitude. Additionally, comparisons of potential and kinetic energy confirm that the combined approach effectively captures the energy exchange during rogue wave formation across all tested intervals.
@@ 7 Discussion and Conclusion
Regarding our initial motivation to evaluate RC’s performance in predicting unseen data, we have observed a remarkable result: RC can synchronize well with a higher-order maximal intensity breather and random ocean waves, despite these not being part of the training data. Nevertheless, significant improvements, particularly in autonomous prediction, can be achieved by combining quasi-periodic waves with random waves in the training data. This finding suggests that to effectively learn the dynamics of spatio-temporal Hamiltonian systems, the training data must sufficiently sample the phase space, particularly the invariant manifold—a requirement that poses additional challenges for training.
@@ Appendix B Maximum Local Lyapunov Exponent Calculation
For the training data shown in Fig. 2, the corresponding analysis in Fig. 14 shows that the stretching number becomes positive during the onset of instability and negative during compression phases. The amplitude of the stretching number closely tracks the most active unstable mode and remains temporally synchronized with the envelope dynamics. Note that the largest Lyapunov exponent—defined as the long-time average of the stretching number—is not well-defined in this context due to persistent fluctuations over time.

### v2

@@ 4 Maximal Intensity Breather
We use the same reservoir, i.e., RC_AB, as Sec. 3 without retraining. This means that we use the generated $W_{x},W_{u}$ along with the learned $W_{o}$ to generate hidden states $\mathbf{x}^{(k)}$ from the input testing data shown in Fig. 7. We discard $N_{w}$ transient points before calculating one-step predictions, as presented in Fig. 8. The NRMSE of one-step prediction is displayed in panel (a), with the error axis shown on a logarithmic scale. The figure shows that the one-step error is generally insignificant, except near the peak of the breather. Even though the multiple lobes are not present in the RC training data, the RC’s one-step prediction profile is remarkably similar to the expected result, as observed in panels (b) and (c). Furthermore, although the RC prediction slightly underestimates the highest peak, as shown in panel (d), the RC successfully captures the overall shape of the highest peak breather.
Figure 7: A fifth-order maximum intensity breather as testing data.
Figure 8: RC_AB can accurately forecast the dynamics of maximum intensity breather in Fig. 7 using one-step prediction, even though the RC was trained using data in Fig. 2. (a) The NRMSE of RC prediction is shown on a logarithmic scale. (b)-(d) Snapshots comparing RC prediction and maximal intensity breather data at $t=4.75,5.75,6.25$.
Figure 9 shows the prediction horizon of the maximal intensity breather testing data, with a legend similar to Fig. 4. Overall, the prediction horizon in the maximum intensity breather is shorter than in Section 2.3. Despite the shorter prediction horizon, our RC can forecast the occurrence of extreme rogue waves with a lead time of 0.5 time units.
The maximal intensity breather represents a path within the homoclinic orbit. It is important to note that the RC model can provide reasonable one-step predictions, even though the original training data do not directly traverse this trajectory. However, due to the inherent sensitivity of the homoclinic orbit, where any infinitesimal deviation leads to exponential divergence, RC struggles with long-term autonomous prediction. The rapid divergence of trajectories near the homoclinic orbit highlights the challenge of maintaining accurate forecasts beyond short horizons.
Figure 9: RC can do a short time forecasting of propagation of maximal intensity breather. The blue dots represent prediction horizons. The red line represents the maximum of the wave envelope.
@@ 5 Random Ocean Wave Simulation
This section demonstrates the forecasting performance of RC in predicting the occurrence of rogue waves in random ocean waves. To that end, we synthesize long-crested irregular waves in deep water. The initial condition for the simulation follows the JONSWAP spectrum, with the detailed procedure given in Appendix C. The physical conditions correspond to deep water waves with a peak period $T_{p}=8$ s and a significant wave height $H_{s}=8$ m. According to the linear ocean wave dispersion relation, this corresponds to a water depth of at least 50 mand a significant wave steepness of $H_{s}/L_{p}=0.08$, where $L_{p}$ is the wavelength at the peak period. We then generate several sets of data with different random phase realizations.
@@
Figure 10: RC can capture the location of rogue waves in random ocean waves for one-step prediction. The plot shows the envelope profile at three distinct times $t=1.92,4.78,9.63$.
Figure 11: Autonomous prediction for random ocean waves. The bottom curve shows the maximal envelope amplitude with dimensionless time $t$ (bottom axis) and corresponding physical time $T$ (top axis). $t_{0}$ and $T_{0}$ are the start of forecasting time in dimensionless and physical time respectively. The upper figure corresponds to the profile of autonomous prediction at times corresponding to (a), (b), and (c).
For autonomous prediction, we initiate the forecasting process at $t_{0}$ which is 0.13 dimensionless time units prior to the rogue wave event, which corresponds to 1.32 s in physical time. The transformation between dimensionless time $t$ and physical time $T$ is given in Eq. (18). The results, presented in 11, show that both RC_Cm and RC_AB exhibit reasonable agreement with the NLS simulation data in predicting the occurrence and location of rogue waves. Notably, RC_Cm yields a more accurate estimate of the wave envelope’s amplitude than RC_AB.
@@ 6 Enhancing Autonomous Prediction
$\tilde{u}^{(k)}_{j}=(1-\delta_{j})u_{j}^{(k)}+\delta_{j}\hat{u}_{j}^{(k)},$ (12)
@@
Figure 12: Heatmap of rogue wave dynamics. The upper left panel shows the ground-truth data. The remaining panels shows the autonomous RC_AB prediction with modified procedures: partial update (upper right), normalization of the solution norm (lower left), and a combination of partial update and normalization of solution norm (lower right). The update and normalization procedures accurately capture both the amplitude and timing of the rogue wave.
@@
Overall, all approaches successfully predict the occurrence of rogue waves. However, each method exhibits distinct limitations as clearly evident in Fig 12. The update-only approach tends to underestimate the rogue wave amplitude, while the normalization approach predicts the rogue wave with a slight delay. In contrast, the combined approach accurately captures both the amplitude and timing of the rogue wave.
@@
Next, we examine the sensitivity of the update interval in the combined approach by considering three assimilation intervals: 30, 70, and 90 time steps. The top left panel of the Fig. 12 shows the peak profile of the forecasted rogue wave, demonstrating close agreement in both shape and amplitude. The remaining panels display comparisons of potential and kinetic energy, which confirm that the combined approach effectively captures the energy exchange during rogue wave formation across all tested intervals.
@@ 7 Discussion and Conclusion
Regarding our initial motivation to evaluate RC’s performance in predicting unseen data, we have observed a remarkable result: RC can synchronize well with a higher-order maximal intensity breather and random ocean waves, despite these not being part of the training data. Nevertheless, significant improvements, particularly in autonomous prediction, can be achieved by combining quasi-periodic waves with random waves in the training data. This finding suggests that to effectively learn the dynamics of spatio-temporal Hamiltonian systems, the training data must sufficiently sample the phase space, particularly the invariant manifold, a requirement that poses additional challenges for training.
@@ Appendix B Maximum Local Lyapunov Exponent Calculation
For the training data shown in Fig. 2, the corresponding analysis in Fig. 14 shows that the stretching number becomes positive during the onset of instability and negative during compression phases. The amplitude of the stretching number closely tracks the most active unstable mode and remains temporally synchronized with the envelope dynamics. Note that the largest Lyapunov exponent, defined as the long-time average of the stretching number, is not well-defined in this context due to persistent fluctuations over time.